Arthur is 16years old and his sister Renie 10years old. how many years go was Arthur twice the age of Renie
Answer by sachi(548)

Arthur is 16years old and his sister Renie 10years old
say x yrs ago Arthur was twice the age of Renie
so 16-x=2(10-x)=20-2x
or x=20-16=4
ans